Find the quotient and remainder of f(x) = x4 -2 divided by p(x) =x -1.
Please show how to do this.
Thank you

Answer by pradhyumna.kumar@gmail.com(62)

Find the quotient and remainder of f(x) = x4 -2 divided by p(x) =x -1.
ans 4x-2/ x-1 quotient is 4 , remainder is 2 .